# Avoiding strict saddle points of nonconvex regularized problems

### v2

@@ 2.1 Optimality condition
    $\partial f(\bar{x})=\{v\mid f(\bar{x})+\langle v,x-\bar{x}\rangle+o(\|x-\bar{x}\|)\leq f(x),\ {\rm as}\ x\rightarrow\bar{x}\}.$ (2)
@@
Let $s(\cdot)=r\circ|\cdot|$. It is trivial to see $s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$.
The subgradients of $s$ are given in the following proposition.
@@
Obviously, $\hat{s}(\cdot)$ is smooth on $\mathbb{R}$ with $\hat{s}^{\prime}(0)=r^{\prime}(|0^{+}|)>0$ and $s(x)=\max\{\hat{s}(x),-\hat{s}(x)\}$. By [37, Exercise 8.31], we have $\partial s(x_{i})=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$. On the other hand, $\liminf\limits_{x\rightarrow 0,x\neq 0}\frac{s(x)-s(0)-v(x-0)}{|x-0|}\geq 0$ hold true for $v=s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $v=s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$, meaning $r^{\prime}(|0^{+}|),-r^{\prime}(|0^{+}|)\in\hat{\partial}s(0)$. It follows from [37, Theorem 8.6] that $\partial s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]\subset\hat{\partial}s(0)\subset\partial s(0)$. Hence, $\partial s(0)=\hat{\partial}s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$.
@@
Overall, we have shown that (4) is true. It then follows from [37, Proposition 10.5] that (3) is true. ∎
We can appeal to the classical rule of Fermat [37, Theorem 10.1] to characterize a local minimum. If $x^{*}$ is a critical point of (P), then
@@
Notice that if $r^{\prime}(|0^{+}|)=\infty$, (7) holds naturally true. Any point satisfy the necessary optimality conditions (6) and (7) are referred as the critical point.
@@
    $d^{T}\nabla^{2}f(x^{*})d-\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}\geq$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (8)
@@
    (sufficient) A critical point $x^{*}$ of $F$ is a strict local minimum if
    $d^{T}\nabla^{2}f(x^{*})d-\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}>$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (9)
@@
(i) is proved in [38, Theorem 3.3]. We only prove (ii).
@@
It is easy to know that $x^{*}$ is a strict local minimum of $F_{|\mathcal{I}^{*}}(x)$ since $x^{*}$ is the critical point of (P) and (8) holds. Therefore for sufficiently large $k$,
@@
2.2 Strict Saddle Point
@@
A critical point $x^{*}$ is a strict saddle point of $F(x)$ if $d^{T}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})d<0$ for $\forall d\in\mathcal{M}_{\mathcal{I}^{*}}\setminus\{0\}$. If all critical points of $F(x)$ are either strict saddle points or local minimums, $F$ satisfies the strict saddle property.
This definition is related to the “active strict saddle point” proposed in [23], in which it is shown that the presence of negative curvature alone in nonsmooth settings does not guarantee escape from saddle points. To address this, they define the active strict saddle point for nonsmooth problems as the saddle point on the active manifold containing the critical point on which the function is partly smooth. Our definition also belongs to this kind, as $\mathcal{M}_{\mathcal{I}^{*}}$ is an active manifold containing $x^{*}$.
2.3 Center Stable Manifold Theorem
As many other works in smooth setting like [40], we interpret our algorithm as the fixed point iteration of a nonlinear map $T:\mathbb{R}^{2n}\rightarrow\mathbb{R}^{2n}$. Likewise, we define the fixed points with at least one eigenvalue of $DT$ larger than $1$ in magnitude as the unstable fixed points, as described below.
@@ 2.3 Center Stable Manifold Theorem
(Unstable fixed point) Let $x^{*},\epsilon^{*}$ be a fixed point of a $C^{1}$ map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. $x^{*},\epsilon^{*}$ is called an unstable fixed point of $T$ if $DT(x^{*},\epsilon^{*})$ has at least one eigenvalue of magnitude strictly larger than 1.
@@
(The Center Stable Manifold Theorem) Let 00 be a fixed point for the $C^{1}$ local diffeomorphism $T:U\rightarrow\mathbb{R}^{n}$ where $U$ is a neighborhood of zero in $\mathbb{R}^{n}$. Let $E^{s}\oplus E^{c}\oplus E^{u}$ be the invariant splitting of $\mathbb{R}^{n}$ into the generalized eigenspaces of $DT(0)$ corresponding to eigenvalues of absolute value less than one, equal to one, and greater than one. There exists a $C^{1}$ embedded disk $W_{loc}^{cs}$ that is tangent to $E_{s}\oplus E_{c}$ at 0 and a neighborhood $B$ around 0 such that $T(W_{loc}^{cs})\cap B\subset W_{loc}^{cs}$, and $\cap_{k=0}^{\infty}T^{-k}(B)\subset W_{loc}^{cs}$.
The CSM theorem only guarantees the avoidance of strict saddle point in the neighborhood for local diffeomorphism, and this local result needs to be extended to the global setting to guarantee that the set of almost all initial points that converge to any unstable fixed point has measure zero. Such result was established in [40, Theorem 2] for a global diffeomorphism map, under which the preimage of a measure zero set also has measure zero. The following corollary in [23] shows the same conclusion for lipeomorphism (A map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is called a lipeomorphism if it is a bijection and both $T$ and $T^{-1}$ are Lipschitz continuous).
@@
[23, Corollary 12] Let $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a lipeomorphism and let $\mathcal{U}_{T}$ consists of all unstable fixed points $x$ of $T$ at which the Jacobian $\nabla T(x)$ is invertible. Then, the set of initial points attracted by such fixed points
$W:=\{x\in\mathbb{R}^{n}:\lim_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$
has zero Lebesgue measure.
@@ 3 Damped Iterative Reweighted $\ell_{1}$ Algorithms
where $w(x_{i})=r^{\prime}(|x_{i}|)$. To leverage the CSM theorem, the map $S$ needs to be a local diffeomorphism at a critical point and a Lipeomorphism globally. However, this is not true in our case since $S$ is generally non-invertible. To see this, one can rewrite the map $S$ using the soft-threshold operator with a changing threshold as
$S(x^{k})=\mathbb{S}_{\lambda w^{k}}(x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})).$
@@
$\min\limits_{\theta\in\mathbb{R}^{2n}}F(\theta):=f(x)+\lambda\sum\limits_{i=1}^{n}r(|x_{i}|+\epsilon_{i}^{2}).$ (13)
with $\theta:=[x;\epsilon]$ for shorthand. Here we also treat $\epsilon$ as another variable to optimize, so that the algorithm updates both $x$ and $\epsilon$ by solving the following subproblem
$x^{k+1}=S(x^{k},\epsilon^{k})=\arg\min_{y}G(y;x^{k},\epsilon^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w_{i}^{k}|y_{i}|,$
with $w_{i}^{k}=w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+\epsilon_{i}^{k})$ and setting $\epsilon^{k+1}=\mu\epsilon^{k}$ with $\mu\in(0,1)$.
The subproblem solution can be given explicitly using the solft-thresholding operator. This update can be written as
$S(x^{k})=\mathbb{S}_{\lambda w^{k}/\beta}(x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})),$ (14)
@@
$S_{i}(x^{k})=\left\{\begin{aligned} &x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i}),&\text{ if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i})>0,\\ &x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i}),&\text{ if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i})<0,\\ &0,&\text{otherwise }.\end{aligned}\right.$ (15)
The map $S$ is not a bijection, as it maps multiple values to 00. To address this issue, we need to revise this algorithm to ensure it satisfy these conditions. Motivated by the technique used in [23], we modified the algorithm with following damping iteration,
$x^{k+1}=(1-\alpha)x^{k}+\alpha S(x^{k},\epsilon^{k}),$
where $\alpha\in(0,1)$ is the damping parameter. Intuitively, if $S_{i}(x^{k},\epsilon^{k}_{i})=0$, the damping operator maps $x^{k}_{i}$ to $(1-\alpha)x^{k}_{i}$, resulting in a bijective map. We can use the iterate $(x;\epsilon)\in\mathbb{R}^{2n}$ and equivalently write the update of $(x^{k};\epsilon^{k})$ as
$(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})=\begin{bmatrix}(1-\alpha)x^{k}+\alpha S(x^{k},\epsilon^{k})\\ \mu\epsilon^{k}\end{bmatrix}$ (16)
with $\epsilon_{0}>0$, $\mu\in(0,1)$ and $\alpha\in(0,1)$.
@@
$x^{*}$ is a critical point of (P) if and only if $x^{*}=S(x^{*},0)$. In addition, $x^{*}$ is a critical point of (P) if and only if $(x^{*},\epsilon^{*})=T(x^{*},\epsilon^{*})$.
@@
Suppose $x^{*}$ is a critical point. It follows from (6) that if $x_{i}^{*}>0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)]=x_{i}^{*}>0$, implying $S_{i}(x^{*},0)=x_{i}^{*}$. This result also holds true by the same argument if $x_{i}^{*}<0$.
As for $x_{i}^{*}=0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|0^{+}|)]<0$ by (7). Same argument also yields $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|0^{+}|)]>0$. Therefore, $S_{i}(x^{*},0)=0=x_{i}^{*}$. Overall, we have shown that $x^{*}=S(x^{*},0)$ in all cases.
Conversely, suppose $x^{*}=S(x^{*},0)$. If $x_{i}^{*}>0$, it must be true that $x_{i}^{*}=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]>0$, implying $0=x_{i}^{*}-S_{i}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]$. Same argument yields $0=x_{i}^{*}-S_{i}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|x_{i}^{*}|)]$ if $x_{i}^{*}<0$. Therefore, $\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)=0$ for $i\in\mathcal{I}^{*}$.
If $x_{i}^{*}=0$, it indicates that $|\frac{1}{\beta}\nabla_{i}f(x^{*})|=|x_{i}^{*}-\frac{1}{\beta}\nabla_{i}f(x^{*})|\leq\tfrac{\lambda}{\beta}w(0,0)=\tfrac{\lambda}{\beta}r^{\prime}(|0^{+}|)$. Hence, $\nabla_{i}f(x^{*})\in[-\lambda r^{\prime}(|0^{+}|),-\lambda r^{\prime}(|0^{+}|)]$. Overall, $x^{*}$ satisfies (5) and is a critical point of (P). By Assumption 2, it satisfies (6) and (7).
@@
4:     Reweighting: $w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+(\epsilon_{i}^{k})^{2})$.
@@
4 Convergence analysis
The analysis focuses solely on the algorithm involving the update of $\epsilon$. It is noteworthy the same conclusion for Algorithm 10 can be readily deduced by setting $\epsilon$ to zero. We first make the following assumptions for the algorithm.

### v3

@@ 2.1 Optimality conditions
    ^$\hat{\partial}f(\bar{x})=\{v\mid f(\bar{x})+\langle v,x-\bar{x}\rangle+o(\|x-\bar{x}\|)\leq f(x),\ {\rm as}\ x\rightarrow\bar{x}\}.$ (2)
@@
It is trivial to see $s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$. The subgradients of $s$ are given in the following proposition.
@@
Obviously, $\hat{s}(\cdot)$ is smooth on $\mathbb{R}$ with $\hat{s}^{\prime}(0)=r^{\prime}(|0^{+}|)>0$ and $s(x)=\max\{\hat{s}(x),-\hat{s}(x)\}$. By [42, Exercise 8.31], we have $\partial s(x_{i})=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$. On the other hand, $\liminf\limits_{x\rightarrow 0,x\neq 0}\frac{s(x)-s(0)-v(x-0)}{|x-0|}\geq 0$ hold true for $v=s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $v=s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$, meaning $r^{\prime}(|0^{+}|),-r^{\prime}(|0^{+}|)\in\hat{\partial}s(0)$. It follows from [42, Theorem 8.6] that $\partial s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]\subset\hat{\partial}s(0)\subset\partial s(0)$. Hence, $\partial s(0)=\hat{\partial}s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$.
@@
Overall, we have shown that (4) is true. It then follows from [42, Proposition 10.5] that (3) is true. ∎
We can appeal to the classical rule of Fermat [42, Theorem 10.1] to characterize a local minimum. If $x^{*}$ is a stationary point of (P), then
@@
Notice that if $r^{\prime}(|0^{+}|)=\infty$, (7) holds naturally true. Any point satisfy the necessary optimality conditions (6) and (7) are referred as the stationary point.
@@
    $d^{T}\nabla^{2}f(x^{*})d+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}\geq$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (8)
@@
    (sufficient) A stationary point $x^{*}$ of $F$ is a strict local minimum if
    $d^{T}\nabla^{2}f(x^{*})d+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}>$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (9)
@@
(i) is proved in [43, Theorem 3.3]. We only prove (ii).
@@
It is easy to know that $x^{*}$ is a strict local minimum of $F_{|\mathcal{I}^{*}}(x)$ since $x^{*}$ is the stationary point of (P) and (8) holds. Therefore for sufficiently large $k$,
@@
2.2 Strict Saddle Points
@@
A stationary point $x^{*}$ is a strict saddle point of $F(x)$ if $d^{T}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})d<0$ for $\forall d\in\mathbb{R}^{|\mathcal{I}^{*}|}\setminus\{0\}$. If all stationary points of $F(x)$ are either strict saddle points or local minimums, $F$ satisfies the strict saddle property.
This definition is related to the “active strict saddle point” proposed in [23], in which it is shown that the presence of negative curvature alone in nonsmooth settings does not guarantee escape from saddle points. To address this, they define the active strict saddle point for nonsmooth problems as the saddle point on the active manifold containing the stationary point on which the function is partly smooth. Our definition also belongs to this kind, as $\mathcal{M}_{\mathcal{I}^{*}}$ is an active manifold containing $x^{*}$.
To show that the strict saddle property is mathematical generic, many works show that this property holds true for a full measure of linear perturbation [23, Theorem 2.9], [44, Theorem 2]. Namely, the perturbed function has the strict saddle property for almost all $v\in\mathbb{R}^{n}$. In our case, we also have the same results by considering the following perturbed nonsmooth regularization problem for $v\in\mathbb{R}^{n}$,
$\min\limits_{x\in\mathbb{R}^{n}}F_{v}(x):=F(x)-\langle v,x\rangle=f(x)+\lambda\psi(x)-\langle v,x\rangle.$ (13)
To show the desired result, we need the assistance of the famous Sard’s theorem. Our proof of the strict saddle property mainly follows the proof of [45, Lemma 2.7] which is proved firstly in [46].
@@
Let $f:\mathcal{M}\rightarrow\mathcal{N}$ be a smooth map of manifolds, if the derivative at $x$ is singular (Jacobian is not full rank), then $x$ is called a critical point of $f$.
Theorem 5.
(Sard’s Theorem) Let $f:\mathcal{M}\rightarrow\mathcal{N}$ be a smooth map of manifolds, and let $C$ be the set of stationary points of $f$ in $\mathcal{M}$. Then $f(C)$ has measure zero in $\mathcal{N}$.
The genericity of the strict saddle property is stated below.
Theorem 6.
For a full Lebesgue measure set of perturbations $v\in\mathbb{R}^{n}$, the perturbation function $F_{v}(x):=F(x)-\langle v,x\rangle$ satisfies
$\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d\neq 0,$ (14)
for all stationary points $x$ of $F_{v}$ with $\mathcal{I}=\mathcal{I}(x)$.
Proof.
Let $s\in\{+1,-1,0\}^{n}$ be a vector of sign pattern and
$\mathcal{M}(s)=\{x\in\mathbb{R}^{n}\mid\text{sign}(x)=s\}$
be a manifold. Let $v:\mathcal{M}(s)\rightarrow\mathbb{R}^{n}$ be a smooth map: $[v(x)]_{\mathcal{I}}=\nabla_{\mathcal{I}}F(x)$ and $[v(x)]_{\mathcal{I}^{c}}=x$ with $\mathcal{I}=\mathcal{I}(s)$. The Jacobian of $v$ at $x$ is then given by $\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)$. By Sard’s Theorem,
$\{[\nabla_{\mathcal{I}}F(x);v_{\mathcal{I}^{c}}]\mid\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)\text{ is singular},x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n}$
$\implies$ $\{v\mid v=v(x),\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)\text{ is singular},x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n}$
$\implies$ $\mathcal{V}(s):=\{v\mid 0\in\partial F_{v}(x),\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d=0,x\in\mathcal{M}\}\text{ has measure 0 in }\mathbb{R}^{n}.$
Now we consider all the possible sign pattern vectors $s\in\{+1,-1,0\}^{n}$. There are at most $3^{n}$ different patterns. Therefore,
$\bigcup_{s\in\{+1,-1,0\}^{n}}\mathcal{V}(s)=\{v\mid 0\in\partial F_{v}(x),\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d=0,x\in\mathbb{R}^{n}\}$
has measure 0 in $\mathbb{R}^{n}$. ∎
Center Stable Manifold Theorem
As many other works in smooth setting like [47], we interpret our algorithm as the fixed point iteration of a nonlinear map $T$. Likewise, we define the fixed points with at least one eigenvalue of $DT$ larger than $1$ in magnitude as the unstable fixed points, as described below.
Definition 5.
(Unstable fixed point) Let $x^{*}$ be a fixed point of a $C^{1}$ map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. $x^{*}$ is called an unstable fixed point of $T$ if $DT(x^{*})$ has at least one eigenvalue of magnitude strictly larger than 1.
@@
(The Center Stable Manifold Theorem) Let 00 be a fixed point for the $C^{1}$ local diffeomorphism $T:U\rightarrow\mathbb{R}^{n}$ where $U$ is a neighborhood of zero in $\mathbb{R}^{n}$. Let $E^{s}\oplus E^{c}\oplus E^{u}$ be the invariant splitting of $\mathbb{R}^{n}$ into the generalized eigenspaces of $DT(0)$ corresponding to eigenvalues of absolute value less than one, equal to one, and greater than one. There exists a $C^{1}$ embedded disk $W_{loc}^{cs}$ that is tangent to $E_{s}\oplus E_{c}$ at 0 and a neighborhood $B$ around 0 such that $T(W_{loc}^{cs})\cap B\subset W_{loc}^{cs}$. In addition, if $T^{k}(x)\in B$ for all $k\geq 0$, then $x\in W_{loc}^{cs}$.
If $\nabla T(0)$ is invertible, then $T$ is locally a diffeomorphism around 0. The CSM theorem only guarantees the avoidance of strict saddle point in the neighborhood for local diffeomorphism, and this local result needs to be extended to the global setting to guarantee that the set of almost all initial points that converge to any unstable fixed point has measure zero. This requires the map has the so-called 0-property, meaning the preimage of a measure zero set under this map still has measure zero. Such result was established in [47, Theorem 2] for a global diffeomorphism map and then for a globally lipeomorphism (A map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is called a lipeomorphism if it is a bijection and both $T$ and $T^{-1}$ are Lipschitz continuous) in [23]. Moreover, to extend this property from a particular stationary point $x^{*}$ to the set of all stationary points, this can be shown by using the fact that $\mathbb{R}^{n}$ is second countable no matter the number of stationary points are countable (see [23, Corollary 2.12] for a standard proof). We list their results below.
Theorem 8.
[47, Theorem 2] Let $g$ be a $C^{1}$ mapping from $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $det(DT(x))\neq 0$ for all $x\in\mathbb{R}^{n}$. Let $\mathcal{U}_{T}$ consists of all unstable fixed points $x$ of $T$ Then the set of initial points that converge to an unstable fixed point has measure zero, $\mathscr{U}=\{x\in\mathbb{R}^{n}:\lim_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$ has zero Lebesgue measure.
@@
[23, Corollary 12] Let $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a lipeomorphism and let $\mathcal{U}_{T}$ consists of all unstable fixed points $x$ of $T$ at which the Jacobian $\nabla T(x)$ is invertible. Then, the set of initial points attracted by such fixed points $\mathscr{U}=\{x\in\mathbb{R}^{n}:\lim\limits_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$ has zero Lebesgue measure.
To summarize, the analysis of escaping from strict saddle points under the strict saddle property generally consists of three steps:
  1. (1).
    Interpret the iterative algorithm as a fixed point iteration, and build the equivalence of the unstable fixed points and the strict saddle points.
  2. (2).
    Show the fixed point iteration is locally a $C^{1}$ local diffeomorphism around a strict saddle point (e.g., a $C^{1}$ smooth map with invertible Jacobian at fixed point), so that the CSM theorem guarantees the avoidance of converging to the strict saddle point.
  3. (3a).
    Show the fixed point iteration is globally a diffeomorphism.
  4. (3b).
    (Alternative to (3a)) Show the fixed point iteration is globally a lipeomorphism.
@@ 3 Damped Iterative Reweighted $\ell_{1}$ Algorithms
where $w(x_{i})=r^{\prime}(|x_{i}|)$. To leverage the CSM theorem, the map $S$ needs to be a local diffeomorphism at a stationary point and a Lipeomorphism globally. However, this is not true in our case since $S$ is generally non-invertible.
@@
$\min\limits_{x,\epsilon\in\mathbb{R}^{n}}F(x,\epsilon):=f(x)+\lambda\sum\limits_{i=1}^{n}r(|x_{i}|+\epsilon_{i}).$ (15)
Here we also treat $\epsilon$ as variables, so that the algorithm updates $x$ and $\epsilon$ by $(x^{k+1},\epsilon^{k+1})=S(x^{k},\epsilon^{k})$
$\begin{cases}x^{k+1}=S^{x}(x^{k},\epsilon^{k})=\arg\min\limits_{y}G(y;x^{k},\epsilon^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum\limits_{i=1}^{n}w_{i}^{k}|y_{i}|,\\ \epsilon^{k+1}=S^{\epsilon}(x^{k},\epsilon^{k})=\mu\epsilon^{k}\end{cases}$
with $w_{i}^{k}=w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+\epsilon_{i}^{k})$ and $\mu\in(0,1)$.
The subproblem solution can be given explicitly using the solft-thresholding operator,
$S^{x}(x^{k},\epsilon)=\mathbb{S}_{\lambda w^{k}/\beta}(x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})),$ (16)
@@
$S^{x}_{i}(x^{k},\epsilon^{k})=\begin{cases}x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i}),&\text{if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i})>0,\\ x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i}),&\text{if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i})<0,\\ 0,&\text{otherwise }.\end{cases}$ (17)
$S$ is not a bijection, as it maps multiple values to 00. Motivated by the technique used in [23], we propose to use the damped operator: $T=(1-\alpha)I+\alpha S$, which can be written explicitly as
$(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})=\begin{bmatrix}(1-\alpha)x^{k}+\alpha S^{x}(x^{k},\epsilon^{k})\\ (1-\alpha(1-\mu))\epsilon^{k}\end{bmatrix}$ (18)
with $\epsilon_{0}\in\mathbb{R}^{n}_{++}$ and $\alpha\in(0,1)$.
@@
$x^{*}$ is a stationary point of (P) if and only if $x^{*}=S^{x}(x^{*},0)$. In addition, $x^{*}$ is a stationary point of (P) if and only if $(x^{*},\epsilon^{*})=T(x^{*},\epsilon^{*})$.
@@
Suppose $x^{*}$ is a stationary point. It follows from (6) that if $x_{i}^{*}>0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)]=x_{i}^{*}>0$, implying $S_{i}^{x}(x^{*},0)=x_{i}^{*}$. This result also holds true by the same argument if $x_{i}^{*}<0$.
As for $x_{i}^{*}=0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|0^{+}|)]<0$ by (7). Same argument also yields $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|0^{+}|)]>0$. Therefore, $S_{i}^{x}(x^{*},0)=0=x_{i}^{*}$. Overall, we have shown that $x^{*}=S^{x}(x^{*},0)$ in all cases.
Conversely, suppose $x^{*}=S^{x}(x^{*},0)$. If $x_{i}^{*}>0$, it must be true that $x_{i}^{*}=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]>0$, implying $0=x_{i}^{*}-S_{i}^{x}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]$. Same argument yields $0=x_{i}^{*}-S_{i}^{x}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|x_{i}^{*}|)]$ if $x_{i}^{*}<0$. Therefore, $\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)=0$ for $i\in\mathcal{I}^{*}$.
If $x_{i}^{*}=0$, it indicates that $|\frac{1}{\beta}\nabla_{i}f(x^{*})|=|x_{i}^{*}-\frac{1}{\beta}\nabla_{i}f(x^{*})|\leq\tfrac{\lambda}{\beta}w(0,0)=\tfrac{\lambda}{\beta}r^{\prime}(|0^{+}|)$. Hence, $\nabla_{i}f(x^{*})\in[-\lambda r^{\prime}(|0^{+}|),-\lambda r^{\prime}(|0^{+}|)]$. Overall, $x^{*}$ satisfies (5) and is a stationary point of (P). By Assumption 2, it satisfies (6) and (7).
@@
4:     Reweighting: $w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+\epsilon_{i}^{k})$.
@@
3.1 Convergence analysis
It is noteworthy that the same conclusion for Algorithm 10 can be readily deduced by setting $\epsilon$ to zero. We first make the following assumptions for the algorithm.